# Data Assimilation with Machine Learning Surrogate Models: A Case Study with FourCastNet

### v1

@@ 7 Conclusions
We empirically show and theoretically justify that filter stability, particularly using 3DVar, is achievable in settings with surrogate weather dynamics models, which allow for substantial computational speedup compared to filtering with NWP models. We additionally show that forecasting with 3DVar initializations produced from a data-driven weather surrogate can provide more accurate short-term predictions than with more naive approaches to initialization, such as using interpolated observations. We note that results using 3DVar with data-driven weather surrogates can potentially be improved with a more physically-informed choice of background covariance $C$ by providing improved stability.
The success of our filtering experiments in our global setting offers promise to future directions. Specifically, a more challenging yet important future direction is assimilating real observations, which involves data collected irregularly over the globe, and measurements that may not directly correspond to an atmospheric feature in the latent space. In these settings, a nonlinear observation operator is needed to transform the observations from quantities that indirectly measure the feature of interest into that feature. Future research in this direction can assess issues that arise due to non-linearities in the observation process, as well as the impact of associated measurement noise on the long-term filtering stability.
We emphasize that, despite the fact that our observational dataset (1) lies on a regular grid, (2) directly measures the quantities in our latent space, and (3) has low measurement error, our results provide positive implications for the task of assimilating coarse NWP forecasts. Assimilating these coarse forecasts can be done to reduce the cost of expensive NWP solvers while still providing high-resolution, accurate estimates of the latent space. These coarse NWP forecasts can provide important information to assimilate into surrogate models’ forecasts, particularly in maintaining longer horizon forecast accuracy, at a cost substantially cheaper than high-resolution NWP solves. Since our results show physically realistic-looking filtering results that have stable errors over a long time horizon, data assimilation with weather surrogates shows substantial promise for applications in this direction.
@@ 8 Acknowledgments
Appendix B Proof of Theorem 1
@@
$x_{t}^{o}$ $=(I-KH)\mathcal{F}(x_{t-1}^{o})+\underbrace{KH\mathcal{F}(x_{t-1}^{\text{true}})+\gamma K\eta_{t}}_{=KH\mathcal{F}(x_{t-1}^{o})},$ (13)
@@
$x_{t}^{s}$ $=(I-KH)\mathcal{F}_{s}(x_{t-1}^{s})+\underbrace{KH\mathcal{F}_{s}(x_{t-1}^{\text{true}})+\gamma K\eta_{t}}_{=KH\mathcal{F}_{s}(x_{t-1}^{s})}$
$x_{t}^{o}$ $=(I-KH)\mathcal{F}(x_{t-1}^{o})+\underbrace{KH\mathcal{F}(x_{t-1}^{\text{true}})+\gamma K\eta_{t}}_{=KH\mathcal{F}(x_{t-1}^{o})},$
@@
$=(I-KH)\left[\mathcal{F}_{s}(x_{t-1}^{s})-\mathcal{F}(x_{t-1}^{s})\right]+(I-KH)\left[\mathcal{F}(x_{t-1}^{s})-\mathcal{F}(x_{t-1}^{o})\right]$
[…]
where we utilize the assumption in (8) to upper-bound the first term and the mean-value-theorem for vector-valued functions to rewrite the second term. Taking the norm of both sides gives
@@
where we again use the assumption of inequality (7) in Theorem 1. By taking the expectation of both sides of the inequality, we obtain the inequality
@@
Figure 11: Visualization of three ensemble members’ 48 hour forecasts of 10m wind speed for May 25, 2023 using three different initializations: noisy ground truth ERA5 data, noisy interpolated $4.5^{\circ}$ observations, and a 3DVar analysis ($4.5^{\circ}$ observations). For visual reference, we include ground truth ERA5 10m wind speed on May 25, 2023.
Figure 12: Visualization of three ensemble members’ 48 hour forecasts of mean sea level pressure for May 25, 2023 using three different initializations: noisy ground truth ERA5 data, noisy interpolated $4.5^{\circ}$ observations, and a 3DVar analysis ($4.5^{\circ}$ observations). For visual reference, we include ground truth ERA5 10m wind speed on May 25, 2023.
Appendix C Evaluation metrics
Our results primarily evaluate our results on two metrics: latitude-weighted root mean square error (RMSE) and latitude-weighted anomaly correlation coefficient (ACC). We use these two metrics based on the analysis done in Pathak et al. [2022], which utilized these metrics defined in Rasp et al. [2020].
The latitude-weighted root-mean-square-error (RMSE) at time $t$ across $f$ different features is defined in equation (16) as
@@ Appendix C Evaluation metrics
We additionally utilize the latitude-weighted anomaly correlation coefficient (ACC) at time $t$ across $f$ different features, defined in equation (18) as
@@
Appendix D Assimilation visualizations for all atmospheric features
Figures 9 and 10 visualize the ground truth ERA5 data, interpolated $4.5^{\circ}$ noisy observations, and 3DVar analyses using these $4.5^{\circ}$ observations for all 20 atmospheric features. The 3DVar analyses were constructed from 365 days of assimilating sparse, noisy $4.5^{\circ}$ resolution observations every 6 hours.
The rows of Figure 9 visualize, in order, U-component wind speed at 10 meters above the surface, V-component wind speed at 10 meters above the surface, temperature at 2 meters above the surface, surface pressure, mean sea level pressure, temperature at pressure level 850 hPa, U-component wind speed at pressure level 1000 hPa, V-component wind speed at pressure level 1000 hPa, geopotential at pressure level 1000 hPa, and U-component wind speed at pressure level 850 hPa.
@@ Appendix D Assimilation visualizations for all atmospheric features
Appendix E Sample 48 hour ensemble forecasts for Typhoon Mawar, 2023
To supplement the single ensemble member visualizations in Figures 7a and 7b, we include visualizations for the 48 hour forecasts of three other randomly selected ensemble members for each of our initial conditions in Figures 11 and 12. As mentioned in the main text, the forecasts from the interpolated $4.5^{\circ}$ observations generally tend to result in less extreme predictions compared to the ground truth or the forecasts from initializing with noisy ground truth ERA5 data or our 3DVar analysis (4.5 observations).

### v2

@@ 7 Conclusions
We empirically show and theoretically justify that filter stability, particularly using 3DVar, is achievable in settings with surrogate weather dynamics models, which allow for substantial computational speedup compared to filtering with NWP models. We additionally show that forecasting with 3DVar-based initializations produced from a data-driven weather surrogate can provide more accurate short-term predictions than with more naive approaches to initialization, such as using interpolated observations. We note that results using 3DVar with data-driven weather surrogates can potentially be improved with a more physically-informed choice of background covariance $C$ by providing improved stability.
The success of our filtering experiments in our global setting offers promise to future directions. Specifically, a more challenging yet important future direction is assimilating real observations, which involves data collected irregularly over the globe, and measurements that may not directly correspond to an atmospheric feature of interest. In these settings, a nonlinear observation operator is needed to transform the observations from quantities that indirectly measure the feature of interest into that feature. Future research in this direction can assess issues that arise due to non-linearities in the observation process, as well as the impact of associated measurement noise on the long-term filtering stability.
We emphasize that, despite the fact that our observational dataset (1) lies on a regular grid, (2) directly measures the quantities of interest, and (3) has low measurement error, our results provide positive implications for the task of assimilating coarse NWP forecasts. Assimilating these coarse forecasts can be done to reduce the cost of expensive NWP solvers while still providing high-resolution, accurate estimates of atmospheric states. These coarse NWP forecasts can provide important information to assimilate into surrogate models’ forecasts, particularly in maintaining longer horizon forecast accuracy, at a cost substantially cheaper than high-resolution NWP solves. Since our results show physically realistic-looking filtering results that have stable errors over a long time horizon, data assimilation with weather surrogates shows substantial promise for applications in this direction.
@@ 8 Acknowledgments
9 Data Availability
In this work, we utilized three datasets: (1) the ERA5 reanalysis data, which can be downloaded at the Copernicus Climate Data Store, Hersbach et al. [2023] for ERA5 data on pressure levels and Sabater [2019] for ERA5 data on land, (2) ECMWF’s IFS-HRES forecasts initialized at either 00:00 UTC or 12:00 UTC daily, which can be downloaded at the TIGGE Data Retrieval portal [Bougeault et al., 2010], and (3) NOAA’s International Best Track Archive for Climate Stewardship (IBTrACS) data [Knapp et al., 2010; Gahtan et al., 2024].
@@ 9 Data Availability
Appendix B Filter and Forecasting Divergence
In this appendix, we visualize instances of divergence that we encountered in our experiments.
Filter divergence without forecast smoothing.
Figure 9 shows a 10m U-component wind speed estimated state from 3DVar for assimilating 4.5 observations without applying a smoothing convolution on FourCastNet’s forecasts, meaning $\mathcal{F}_{s}:=\mathcal{F}_{\text{FCN}}$ in equation (6). The analysis state exhibits early-stage filter divergence for this assimilation task near the top right corner of the image. The initial sign of filter divergence appeared on January 5, 2023 at 12:00 UTC, so the filter was only able to assimilate about 4 days of 6-hourly observational data before showing signs of degradation. There is no obvious physical meaning behind the origin of the filter divergence, so it may be reasonable to assume that the 3DVar analysis inputs to FourCastNet are different enough from the ERA5 dataset that instabilities arise, particularly near the northern pole. FourCastNet version 1 is known to degrade near the poles when applied autoregressively [Bonev et al., 2023a], so an interesting future direction would be to assess FourCastNet v2’s stability in data assimilation tasks.
Figure 9: Example visualization of the 3DVar analysis without forecast smoothing (i.e., $\mathcal{F}_{s}:=\mathcal{F}_{\text{FCN}}$) for U-component wind speed 10m above the surface. This 3DVar analysis corresponds to January 5, 2023 12:00 UTC after assimilating 4.5 observations 6-hourly starting January 1, 2023 00:00 UTC. In this visualization, the extreme values that are characteristic of filter divergence are filled in with white pixels. A pixel is characterized as diverging and filled in with white if it was 10% larger than either the minimum or maximum ERA5 value for that same time point.
Filter divergence for assimilating 5 observations.
Here, we further investigate the filter divergence for assimilating 5 observations seen in Figures 1a and 1b. Figure 10 visualizes the 3DVar estimate of the state for 10m wind speed on May 29, 2023, corresponding to roughly two days before filter divergence is clearly seen in the MSE and ACC metrics. As shown in Figure 10, the filter divergence originates at the center of an extreme event, specifically Typhoon Mawar. The filter divergence originating from this extreme event suggests that, given our assumptions on $C$ within 3DVar, 5 observations do not provide enough localized information around the typhoon to stabilize FourCastNet’s predictions, and these unstable predictions are not adequately corrected within 3DVar based on our static constructing of $C$. As an additional note, we see that the 3DVar analysis lacks the expected small-scale structure of the eye of the typhoon: the largest wind speed values occur in multiple modes rather than as a ring around a low wind speed eye, which further evidences the unphysical nature underlying the divergence.
Figure 10: Visualization of the observed beginning of filter divergence for the task of assimilating 5 observations into FourCastNet’s predictions compared to the ground truth ERA5. This particular snapshot corresponds to global 10m wind speed on May 29, 2023 at 06:00 UTC. The zoomed in region in the 3DVar analysis and the ground truth ERA5 data localize Typhoon Mawar, which is the origin of the 3DVar filter divergence. The colorbar maps the 10m wind speed from the values 0 to the maximum seen in ERA5, which corresponds to roughly 30 m s-1. Extreme values in the 3DVar analysis above 30 m s-1 were masked in plotting and mapped to a solid white color. In this particular snapshot, the 3DVar analysis had an estimated maximum wind speed of roughly 50 m s-1, which is almost double the value seen in ERA5.
Forecast divergence for initializing forecasts with 3DVar analyses (4.5 observations).
We encountered two particular instances of forecast divergence when initializing with our 3DVar analyses that assimilate 4.5 observations. In particular, initializing on August 1, 2023 at 12:00 UTC and 18:00 UTC lead to forecast divergence within a 24 hour forecast horizon. Figure 11 shows that when initializing forecasts on August 1, 2023 at 12:00 UTC, the forecast divergence originates at the center of a typhoon in the Pacific Ocean, specifically Typhoon Khanun. The visualization for forecasting with a 3DVar analysis (4.5 obs.) on August 1, 2023 at 18:00 UTC shows similar behavior to Figure 11.
The 3DVar analyses (4.5 obs.) for these two time points notably have estimated wind speeds much larger than is seen in the ERA5 dataset: about 50 m s-1 for the 3DVar analyses and about 30 m s-1 for ERA5. The large overestimate in the initial condition causes instabilities in the forecast that quickly propagate, as is shown in Figure 11.
Figure 11: Visualization of an example 3DVar analysis (4.5 observations) of 10m wind speed that, when used as an initialization for FourCastNet, rapidly leads to a divergent forecast. The left-most image shows the 3DVar analysis (4.5 observations) initialization prior to forecasting with FourCastNet. The middle image shows the predicted global 10m wind speed at a 18 hour forecast horizon, and the right-most image shows the predicted global 10m wind speed at a 42 hour forecast horizon. Extreme predicted values are masked by white pixels. A pixel was determined as extreme if it exceeded the maximum wind speed in the initialization, which is roughly 50 m s-1.
Appendix C Background Error Covariance
We construct our background error covariance with simplified assumptions that lead to computational efficiencies in 3DVar. In particular, we assume that the covariance between two distinct locations is proportional to a truncated Gaussian decay as described in equation (5) and visualized in Figure 12. This construction assumes that the covariance between two distinct locations decreases as the distance between the two locations increases, and when the distance between the two points reaches a certain threshold ($k/2$ in our experiments), the covariance immediately drops to 0. In all four kernels, the scaling parameter $\sigma^{2}$, which controls how quickly the association decays as the distance between two locations increases, is held constant at 8 for each setting. The parameter $\sigma^{2}=8$ was chosen so that the covariance decays gradually within our chosen convolutional kernel size, as shown in Figure 12. The main difference among the four kernels is the distance at which the covariance becomes 0, which is dependent on $k$.
Figure 12: Visualization of the different kernels $W^{(k)}$ from equation (5) for $B$ utilized in our data assimilation tasks for $k=8,10,18,$ and $20$. Each kernel has the shared parameter $\sigma^{2}=8$ in defining the Gaussian decay.
Figure 13: Example visualization demonstrating that $HBB^{\text{T}}H^{\text{T}}$ is a diagonal matrix based on our construction of $B$ and $H$. For simplicity, we show this result on a 1D example. Step 1 shows this example’s choice of $B$ and $H$. The $B$ matrix above reflects a 1D convolutional kernel of length 3 with equal size weights with replication padding, and the choice of $H$ assumes the first and fourth element of the length 5 state vector is observed. Step 2 computes the matrix $BB^{\text{T}}$, which is now a symmetric matrix. The rows and columns with red arrows correspond to the indices that are observed based on our choice of $H$. As an aside, the matrix $BB^{\text{T}}H^{\text{T}}$ can be seen by retaining the columns with arrows shown in $BB^{\text{T}}$. Finally, Step 3 computes $HBB^{\text{T}}H^{\text{T}}$, which is a diagonal matrix. This matrix is constructed from retaining only the elements that have arrows in its row and column, as shown in the $BB^{\text{T}}$ matrix. We note that in the construction of $B$, centered at any particular index, $BB^{T}$ allows for a nonzero correlation of other indices at most two spaces away. However, our choice of $H$ observes every third state, meaning that information from each observed index does not influence the state at another observation index since the two indices have 0 estimated covariance. Therefore, $HBB^{\text{T}}H^{\text{T}}$ can be stored as a vector of size 2 rather than a 2x2 matrix.
In order to maximize computational efficiency, we carefully chose $C$ according to our gridded observations so that $(HCH^{\text{T}}+R)$ is a diagonal matrix, avoiding a matrix inverse that costs $\mathcal{O}(d_{y}^{3})$ and instead computes the inverse of a vector of scalars, an operation with $\mathcal{O}(d_{y})$ computational cost. Since $R$ is already assumed to be diagonal, we only need to construct $C$ to ensure that $HCH^{\text{T}}$ is diagonal. Intuitively, we want to construct $C$ such that any two distinct observation locations have 0 covariance. To ensure this property, we construct the background covariance $C$ based on $BB^{\text{T}}$, where $B$ is a convolution with kernel $W^{(k)}$, where $k=8,10,18,$ and 20 in our experiments. Constructed $C$ via $BB^{\text{T}}$ ensures that no matter our choice of $B$, the resulting $C$ is a proper covariance matrix. The convolution $B$ is applied to a state $X_{t}$ for any time $t$ and applies replication padding at the boundaries, which is naive to the fact that this data is collected over a sphere. Future implementations can implement a padding that reflects that the resulting image lies on a sphere.
Figure 13 visually justifies why our choice of $HCH^{\text{T}}$ is diagonal, allowing for us to avoid a computationally expensive $d_{y}$ matrix inversion. Figure 13 contains a 1D example showing that given a choice of $B$ and $H$ consistent with our experimental setting, $HCH^{\text{T}}$ is diagonal. Since we assume that our observations always lie on a regular grid, and given that our choice of $C$ guarantees that observation information will not propagate to another observation’s location, $HCH^{\text{T}}$ is diagonal.
Appendix D Proof of Theorem 1
@@
$x_{t}^{o}$ $=(I-KH)\mathcal{F}(x_{t-1}^{o})+KH\mathcal{F}(x_{t-1}^{\text{true}})+\gamma K\eta_{t},$ (13)
@@
$x_{t}^{s}$ $=(I-KH)\mathcal{F}_{s}(x_{t-1}^{s})+KH\mathcal{F}_{s}(x_{t-1}^{\text{true}})+\gamma K\eta_{t},$
$x_{t}^{o}$ $=(I-KH)\mathcal{F}(x_{t-1}^{o})+KH\mathcal{F}(x_{t-1}^{\text{true}})+\gamma K\eta_{t},$
@@
$=(I-KH)\left[\mathcal{F}_{s}(x_{t-1}^{s})-\mathcal{F}(x_{t-1}^{s})\right]+(I-KH)\left[\mathcal{F}(x_{t-1}^{s})-\mathcal{F}(x_{t-1}^{o})\right].$
Utilizing the assumption in (8) to upper-bound the first term and the mean-value-theorem for vector-valued functions and the assumption of inequality (7) in Theorem 1 to upper-bound the second term, we obtain that
@@
By taking the expectation of both sides of the inequality, we obtain the inequality
@@
Appendix E Evaluation metrics
We evaluate our results using three metrics: latitude-weighted root mean square error (RMSE), latitude-weighted anomaly correlation coefficient (ACC), and continuous ranked probability score (CRPS). We formulate RMSE and ACC based on the definitions provided in Rasp et al. [2020].
The latitude-weighted RMSE at time $t$ across $f$ different features is defined in equation (16) as
@@ Appendix E Evaluation metrics
We additionally utilize the latitude-weighted ACC at time $t$ across $f$ different features, defined in equation (18) as
@@
We compute the CRPS [Matheson and Winkler, 1976], which is defined as follows,
[…]
where $y$ is an observation, $F$ is the CDF of the forecast distribution, and $\mathds{1}$ is the Heavside function. In our work, the $F$ we provide is an empirical CDF of the forecast distribution based on an ensemble.
Appendix F Assimilation visualizations for all atmospheric features
Figures 14 and 15 visualize the ground truth ERA5 data, interpolated $4.5^{\circ}$ noisy observations, and 3DVar analyses using these $4.5^{\circ}$ observations for all 20 atmospheric features. The 3DVar analyses were constructed from 365 days of assimilating sparse, noisy $4.5^{\circ}$ resolution observations every 6 hours.
The rows of Figure 14 visualize, in order, U-component wind speed at 10m above the surface, V-component wind speed at 10m above the surface, temperature at 2 meters above the surface, surface pressure, mean sea level pressure, temperature at pressure level 850 hPa, U-component wind speed at pressure level 1000 hPa, V-component wind speed at pressure level 1000 hPa, geopotential at pressure level 1000 hPa, and U-component wind speed at pressure level 850 hPa.
@@ Appendix F Assimilation visualizations for all atmospheric features
Appendix G Sample 48 hour ensemble forecasts for Typhoon Mawar, 2023
To supplement the single ensemble member visualizations in Figures 7a and 7b, we include visualizations for the 48 hour forecasts of three other randomly selected ensemble members for each of our initial conditions in Figures 16 and 17. As mentioned in the main text, the forecasts from the interpolated $4.5^{\circ}$ observations generally tend to result in less extreme predictions compared to the ground truth or the forecasts from initializing with noisy ground truth ERA5 data or our 3DVar analysis (4.5 observations).
Figure 16: Visualization of three ensemble members’ 48 hour forecasts of 10m wind speed for May 25, 2023 using three different initializations: noisy ground truth ERA5 data, noisy interpolated $4.5^{\circ}$ observations, and a 3DVar analysis ($4.5^{\circ}$ observations). For visual reference, we include ground truth ERA5 and the IFS-HRES forecast of 10m wind speed on May 25, 2023.
Figure 17: Visualization of three ensemble members’ 48 hour forecasts of mean sea level pressure for May 25, 2023 using three different initializations: noisy ground truth ERA5 data, noisy interpolated $4.5^{\circ}$ observations, and a 3DVar analysis ($4.5^{\circ}$ observations). For visual reference, we include ground truth ERA5 and the IFS-HRES forecast of 10m wind speed on May 25, 2023.